First City Bank pays 8 percent simple interest on its savings account balances, whereas Second City Bank pays 8 percent interest compounded annually. If you made a $65,000 deposit in each bank, how much more money would you earn from your Second City Bank account at the end of 8 years? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

  

Your coin collection contains 45 1952 silver dollars. If your grandparents purchased them for their face value when they were new, how much will your collection be worth when you retire in 2055, assuming they appreciate at an annual rate of 5.6 percent? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

  

You expect to receive $44,000 at graduation in two years. You plan on investing it at 10 percent until you have $179,000. How long will you wait from now? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
Answer #1

1.

amount in first bank = 65,000*8*8% + 65000 = $106600

amount in second bank = 65,000*(1+8%)^8 =$120310.4637

Difference in accounts =$120310.4637-$106600 =$13710.46

2.

Number of years elapsed = 2055 - 1952 = 103

FV = PV *(1+t)^t

Future value = 45*(1+5.6%)^103 = $12319.59

3.

t = log(A/P) / n[log(1 + r)]

A = $179000

P = $44000

r = 0.10

t = 14.72 years

Time to wait from now = 2 years + 14.72 years = 16.72 years

Please leave a like. Thank you.